OpenStudy (anonymous):

Which of the following are measurements of the sides of a right triangle? A.101, 99, 20 B.28, 26, 12 C.17, 14, 6 D.none of the above

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

@satellite73

OpenStudy (freckles):

see if this equation holds a^2+b^2=c^2 where c is the larger of the 3 given to you

OpenStudy (anonymous):

ok so it c

OpenStudy (freckles):

\[6^2+14^2=17^2 \text{ holds ? } \\ \text{ Let's find out } \\ 36+196=289 \\ \] is that equation true is 36+196 equal to 289?